A table clock has its minute hand 4.0 cm long. The average velocity of the tip of the minute hand between 6:00 am to 6:30 pm is :-
(a) 1.1 X 10-4 cm/s
(b) 2.2 X 10-4 cm/s
(c) 3.14 X 10-4 cm/s
(d) 3.3 X 10-4 cm/s

Grade:12th pass

3 Answers

erra akhil
8 years ago
Dear Abhay,
According to your question;
Time taken=12.5hrs=12.5*60*60=45000sec.
θπ + π/24 = 25π/24.
θ = ω t + (α  t) / 2.
Since motion of minutes hands in clock is in constant angular velocity, α = 0.
θ = ω t 
on substituting the values we value of; ω= 0.23 π * 10-4 rad/sec.
we know,
v= r ω
here on calculating v we get;
v = 0.92 π * 10-4 cm/s = 3.14 * 10-4 cm/s.
Approve if my answer helped you.
 
 
erra akhil
8 years ago
Dear Abhay,
                  What is the correct answer? 
 
ASh
6 years ago
anybody answer this question fast!!!!!!! PLS........................................................................................